### v4

# Steering LLMs for Formal Theorem Proving

Abstract
Recent progress in automated theorem proving leverages Large Language Models (LLMs) for their capacity to comprehend informal mathematical statements and generate corresponding formal proofs. Even though these techniques perform well, very little exploration has been done to understand how language models interpret and utilize these informal mathematical cues to generate formal proofs more effectively. To address this, we explore activation steering, a lightweight, inference-time mechanism that identifies linear directions in a model’s residual activations corresponding to informal “thought” traces, and nudges those activations to improve proof construction entirely without finetuning. Unlike previous approaches, activation engineering offers valuable insights into language models’ internal reasoning dynamics encoded in their activation space. We evaluated these activation vectors on two distinct tasks: formal proof generation from formal theorems and formal proof generation from informal problem descriptions. Our contributions are twofold: (1) we purpose an activation-based intervention technique to guide proof synthesis in LLMs; and (2) improve performance across two different decoding strategies without additional training.
Machine Learning, ICML
1 Introduction
Interactive proof assistants such as Lean (de Moura et al.,, 2015), Isabelle (Wenzel et al.,, 2008), and Coq (Barras et al.,, 1999) enable formal verification of mathematical proofs and software by leveraging specialized programming languages (Avigad,, 2023; Ringer et al.,, 2019). Neural theorem proving, which integrates neural models like Large Language Models (LLMs) with these assistants, has emerged as a promising approach to automating formal reasoning and theorem proving (First et al.,, 2023; Polu and Sutskever,, 2020; Polu et al.,, 2022; Yang et al.,, 2023; Welleck,, 2023). This integration is intuitive and powerful: proof assistants enforce formal correctness, while neural models predict and suggest logical steps for proof construction. A key challenge, however, is generalizing proof writing to scenarios unseen during training. Multiple prior studies have demonstrated that informal (natural language, NL) information can enhance theorem proving in language models (Lin et al.,, 2024; Welleck et al.,, 2021, 2022).
Despite these advances, it remains unclear how natural language information influences a model’s reasoning process and whether it can be systematically leveraged to improve proof-writing capabilities. Our work addresses this gap by investigating the mechanistic role of natural language information in LLM reasoning and developing a method to harness it effectively.
We find that the task of generating natural language mathematical context can be encoded as activation vectors in the model’s activation space. Using activation steering (Panickssery et al., 2024b, ; Turner et al.,, 2024; Lucchetti and Guha,, 2024), we capture differences in model behavior driven by natural language inputs, enabling targeted interventions that enhance proof-writing without affecting unrelated aspects of performance. In figure 1 we demonstrate how structured natural language interventions improve the quality and the structure of proof. In comparison to natural sampling leads to counterfeit and incorrect proofs. We evaluate this approach on three open-source LLMs like Lemma (Azerbayev et al.,, 2024), InternLM-2 (Ying et al.,, 2024), and InternLM2.5-StepProver (Wu et al.,, 2024) that are designed for theorem proving through training on mathematics data. Our experiments show that steering vectors with targeted NL information consistently improve proof quality across the MiniF2F and PutnamBench benchmarks.
Our key contributions are:
  1. 1.
    Mechanistic insights into how informal mathematical context and natural language information affects LLM reasoning about formal theorems.
  2. 2.
    Activation vector identification for encoding natural language information to guide formal proof construction in a structured and grounded manner.
  3. 3.
    Activation steering framework yielding consistent gains in proof generation quality on MiniF2F and PutnamBench.
This study not only improves how models reason about specific tasks, but also provides a practical, interpretable tool for steering reasoning. These learning’s provide crucial insights to interpret and understand LLMs for theorem proving better.
2 Related Work
Automatic Theorem Proving and Auto-formalization Prior work on automatic theorem proving typically follows the GPT-f framework (Polu and Sutskever,, 2020), which trains a language model on proof state and next tactic pairs, then proves theorems using this model with best-first tree search. Subsequent work has explored several directions, including data augmentation (Han et al.,, 2022), synthetic data for training language models (Rotella et al.,, 2025; Wang and Deng,, 2020), novel proof search methods (Wang et al.,, 2023), further training through curriculum learning (Polu et al.,, 2022), retrieval augmented generation (Yang et al.,, 2023), and practical tools (Welleck and Saha,, 2023).
Figure 1: In this example we show how steering vectors introduce natural language grounding in the proof improving the quality. We have highlighted the effect of steering vectors in yellow. The lines highlighted cyan represent incorrect tactics leading to incorrect proof.
A parallel thread focusing on improving automated theorem proving has explored auto-formalization (Wu et al.,, 2022), which incorporates informal mathematics into formal theorem proving. This direction recognizes the significant value of connecting informal mathematical intuition with formal proof systems. Draft-Sketch-Prove (Jiang et al.,, 2023) demonstrates that language models can leverage informal proofs to enhance formal proving abilities through few-shot prompting rather than supervised fine-tuning. Similarly, LeanStar (Lin et al.,, 2024) introduces ”thought-augmented reasoning” in each proof step, using informal reasoning to guide tactic prediction. However, approaches like LeanStar typically require supervised fine-tuning of large-language models on synthetic data to improve reasoning capabilities.
Our work addresses the fundamental challenge of understanding the relationship between informal and formal mathematics by learning task of producing thoughts directly in activation space to improve theorem proving at inference time. This approach provides a lightweight alternative to learning informal reasoning patterns through fine-tuning and offers insights into how language models internally represent and connect informal mathematical reasoning with formal proof steps. Understanding this relationship is crucial for developing more effective theorem proving systems that can leverage the complementary strengths of both informal mathematical intuition and formal verification.
Language model representation Our approach is inspired by previous research that shows that it is possible to obtain vector representation of tasks that language model learns in context (Hendel et al.,, 2023; Todd et al.,, 2024). These studies are part of broader body of work that examines the linear representation of features such as truthfulness (Azaria and Mitchell,, 2023; Li et al.,, 2023; Marks and Tegmark,, 2024), sentiment (Tigges et al.,, 2024), harmlessness (Zou et al.,, 2023; Zheng et al.,, 2024), sychophancy (Perez et al.,, 2023; Panickssery et al., 2024a, ; Sharma et al.,, 2024), factual knowledge (Gurnee and Tegmark,, 2024), and refusal (Arditi et al.,, 2024). In addition, recent works have employed sparse autoencoders to identify feature directions in an unsupervised manner (Bricken et al.,, 2023; Huben et al.,, 2024; Templeton et al.,, 2024).
A shared hypothesis across these works is that LLMs represent features or concepts as linear directions in activation space (Mikolov et al.,, 2013; Elhage et al.,, 2021; Nanda et al.,, 2023; Park et al.,, 2024). While recent studies suggest that not all features may be linearly encoded (Engels et al.,, 2025), the linearity assumption has been effective for both concept erasure and model steering. In our work we explore how the concept of generating natural language informal data is a linearly representable concept in activation space which improves the performance of these models to prove formal theorems.
Mechanistic Interpretability and Activation Patching Previous work has focused on localizing and editing factual associations within transformers (Meng et al.,, 2022) and probing hidden representations for high-level knowledge (Li et al.,, 2024; Dong et al.,, 2023). Such studies show implicit evaluations the ability of the model, complementing explicit benchmarks (Dong et al.,, 2023). A key technique in mechanistic interpretability is activation patching (Vig et al.,, 2020; Variengien and Winsor,, 2023), which modifies model activations in the residual stream at the time of inference to influence outputs (Zhao et al.,, 2025). Residual streams is the high-dimensional vector that carries and accumulates information across layers via skip (residual) connections allowing each layer to iteratively refine and add shared representation used for prediction. In figure 2 we demonstrate the process of generating activation vectors on the proof step generation task to enable language models to generate natural language patch that contains context for the next tactic.
For theorem proving, mechanistic interpretability should provide insight onto how well language models can comprehend natural language information and use it to comprehend formal proofs. By dissecting these representations, we can identify failure cases, refine step generation, and enhance the proof generation capability. We observe that effective steering transforms activations to align the model’s responses to a structured format to generate natural language information while writing a proof step that allows them to be better at proofing theorems.
3 Methodology
3.1 Model Choice
We build steering datasets for 7B parameter Llemma (Azerbayev et al.,, 2024), InternLM2 (Ying et al.,, 2024) and InternLM2.5-StepProver (Wu et al.,, 2024). These models are trained to generate formal theorems and on popular programming languages to improve their capability of reason about theorem’s better. InternLM family of models are based on LlaMA architecture (Touvron et al.,, 2023), however Llemma model is trained on CodeLlaMA architecture.
Figure 2: Steering Vectors are computed as difference of means for for $p^{+}$ and $p$
3.2 Tasks and Source Data
We constructed steering pairs from a randomly sampled subset $S$ of the Lean-STaR dataset  [^1] https://huggingface.co/datasets/ScalableMath/Lean-STaR-plus (Lin et al.,, 2024). This subset consists of $10,000$ unique data points. Each of these data points consist of a formal proof step and an informal thought required to generate that step.
Every natural prompt $p\in S$ is a formal proof step of a proof. We then create $p^{+}$ by inserting explicit reasoning step of natural language explanations of the proof state. This process gives us prompts pairs that are required to generate steering vectors.
After generating prompt pairs $(p,p^{+})$, we prompt the model to with both the prompts and collect responses $r_{p}$ and $r_{p^{+}}$. We then filter a subset $s\subset S$ where both $r_{p}$ and $r_{p^{+}}$ are valid proof steps and $r_{p}\neq r_{p^{+}}$. The resulting quadruples $\{p,p^{+},r_{p},r_{p^{+}}\}$ form the basis for steering vector construction.
Formal Theorem to Formal Proof Writing In this task, the model has generate a proof step-by-step for a given formal problem. We use standard benchmark for this task miniF2F (Zheng et al.,, 2022). This benchmark contains 244 test set theorems in Lean. To generate proof for the problem we use best first search as well as beam search. We describe the experiment setup in Section 4.
Informal Problem to Formal Proof Writing This task covers writing proof from informal problem statement. We use PUTNAMBENCH (Tsoukalas et al.,, 2024) as the benchmark for this task. This benchmark includes 657 problems in Lean and 640 problems in Coq. We use both these datasets to evaluate steering vectors.
3.3 Constructing Steering Vectors
To extract robust steering vectors, we first identify the activations where these vectors are linearly represented within the model. We focus on residual stream activations i.e the outputs of each transformer layer. Given an informal thought step we identify the layers where the activation vector is active at time of processing the last token.
To identify the causally relevant layers, we first extract steering vector candidate from every layer using Difference of Means method. This is a widely used technique (Turner et al.,, 2024; Arditi et al.,, 2024). This technique is based on constructing contrastive dataset that differs in a specific concept. For our task the contrastive dataset is $s$ and the concept we want to differentiate is producing NL information with the formal proof-step. We begin with a data set of pairs of prompts and their responses $\{(p_{i},p_{i}^{+},r_{p_{i}},r_{p_{i}^{+}})\}_{i=1}^{N}$, where $p_{i}^{+}$ includes informal reasoning steps and $p_{i}$ is the standard prompt. The model $M$ and extract its residual stream activation at the final token of layer $\ell$, denoted
$v_{p_{i}^{-}}^{\ell}=\mathrm{resid}(M(p_{i}^{-}),\ell)\quad\text{and}\quad v_{p_{i}^{+}}^{\ell}=\mathrm{resid}(M(p_{i}^{+}),\ell).$
To isolate the component capturing natural language information about the proof, we apply the Difference of Means method:
$u^{(\ell)}=\frac{1}{|s_{+}|}\sum_{p^{+}\in S_{+}}v_{p^{+}}^{\ell}\;-\;\frac{1}{|s_{-}|}\sum_{p^{-}\in S_{-}}v_{p^{-}}^{\ell},$
Where $v_{p}^{\ell}$ represents the activation vector of layer $\ell$ for a prompt $p$ and $s_{+}$ represents all the samples in the set $s$ that exhibit the targeted behavior while $s_{-}$ represent all the samples. In this scenario the Difference of Means vector is computed by subtracting overall mean activation from the mean activation of the behavior associated with the example. This isolates the direction in the activation space most associated with the target behavior while reducing the influence of general model biases.
Figure 3: Cosine similarity in InternLM2’s residual stream activations drops in later layers for informal reasoning queries, but gradually recovers.
4 Results and Analysis
Setup We construct proofs using best-first search method which is one of the most popular methods to evaluate the theorem proving ability of a language model (Polu and Sutskever,, 2020; Yang et al.,, 2023; Azerbayev et al.,, 2023; Lin et al.,, 2024). For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$. Following standard practice (Polu and Sutskever,, 2020; Yang et al.,, 2023; Welleck and Saha,, 2023; Lin et al.,, 2024), we assume that the state with maximum negative log-probabilities is the ”best” state. Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average logarithmic probability of each generated token. We expand up to $N$ states and we get a successful proofs search when we reach any proof state with no goals.
We observed that the best first search perform poorly in our case as with the natural language tokens in our response the mean of log probabilities is noisy and inconsistent, making it harder to choose the post promising branch. This observation is consistent to Lean-STaR (Lin et al.,, 2024).
To overcome these limitation we apply Sampling technique proposed by Lean-STaR (Moura and Ullrich,, 2021). Sampling is based on sampling trajectories in parallel for each problem, each time generating at most N tactics. The illegal sampled tactics are ignored during sampling. This means at state $s_{i}$ if the sampled tactic is $a_{i}$ is only added to trajectory $(s_{0},a_{0}),(s_{1},a_{1}),...,(s_{i-1},a_{i-1})$ if $a_{i}$ is a legal tactic otherwise we sample another tactic from the model M.
This technique takes approximately same time on GPU as best-first search but accommodates natural language thoughts.
Figure 4: Effect of patched layers on proof generation in miniF2F-test using InternLM2. Y-axis: verified proofs; X-axis: number of patched layers.
Model Decoding N K S MiniF2F
LLEMMA-7B SEARCH 50 1 32 26.2%
InternLM2-7B SEARCH 50 1 32 30.3%
InternLM2-7B (SFT) SEARCH 50 1 32 30.7%
LEAN-COT (InternLM2-7B) SAMPLING 50 32 1 27.0%
LEAN-COT (InternLM2-7B) SEARCH 50 1 32 25.4%
LEAN-STAR (InternLM2-7B) SAMPLING 50 32 1 29.1%
LEAN-STAR (InternLM2-7B) SEARCH 50 1 32 26.2%
COPRA (WITH GPT-4) CUSTOMIZED - 60 1 29.9%
InternLM-2-Math-Plus SEARCH 100 1 32 43.4%
InternLM-2-StepProver SEARCH 100 1 32 48.8%
InternLM-2.5-StepProver SEARCH 600 2 32 50.7%
DeepSeekProver-V1.5 SEARCH 400 32 16 63.5%
InternLM-2.5-StepProver SEARCH 600 256 32 65.9%
LLEMMA-7B + Activation vector SAMPLING 50 32 1 28.1%
LLEMMA-7B + Activation vector SEARCH 50 1 32 26.3%
InternLM2-7B + Activation vector SAMPLING 50 32 1 32.4%
InternLM2-7B + Activation vector SEARCH 50 1 32 26.8%
InternLM-2.5-StepProver + Activation Vector SEARCH 600 2 32 49.18%
InternLM-2.5-StepProver + Activation Vector SAMPLING 600 2 32 52.4%
InternLM-2.5-StepProver + Activation Vector SEARCH 600 256 32 65.57%
InternLM-2.5-StepProver + Activation Vector SAMPLING 600 256 32 66.39%
Table 1: Pass rates on the miniF2F-test dataset with Lean. S is the number of tactics attempted at each expanded node (assumed to be 1 in sampling), and K is the total number of search or sampling attempts per problem.
Model Sampling Pass at
Problems Solved
Lean4 (Out of 657)
InternLM2 7B SEARCH 4096 4
InternLM2.5-StepProver SEARCH 2x32x600 6
Goedel-Prover-SFT SEARCH 512 7
ABEL SEARCH 596 7
Self-play Theorem Prover SEARCH 3200 8
Kimina-Prover-7B-Distill SEARCH 192 10
InternLM2 7B + Activation Steering SAMPLING 4096 4
InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 7
Coq (out of 412)
InternLM2.5-StepProver SAMPLING 2x32x600 0
InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 1
Table 2: Results on PutnamBench with Lean and Coq.
4.1 Main results
We construct distinct steering vectors for each model variant, enabling us to isolate and analyze the activation-level effects of distilled informal reasoning. A detailed comparison of activation patterns across models is presented in Appendix.
miniF2F Our results on the miniF2F test set are summarized in Table 1. Steering the model’s activations consistently improves performance across decoding strategies when compared to the base model. Notably, we observe substantial gains in pass rates under both best-first search and sampling strategies. One exception is a slight drop in performance for InternLM-2.5-StepProver under best-first search with activation steering. We hypothesize that this is due to the injected natural language content lowering the mean log-probability of tokens, which can affect deterministic decoding strategies more than stochastic ones.
PutnamBench Table 2 reports results on the Lean and Coq subsets of PutnamBench. Our findings further support the effectiveness of activation steering: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Coq problem—an improvement over the base model’s zero success rate. This cross-system transfer underscores the generality and robustness of the distilled reasoning signal encoded in the steering vectors.
4.2 Research Questions
4.2.1 How effective and efficient are activation steering vectors?
Finetuning To understand the quality of our steering vectors, we do Low Rank Adapter fine-tuning (LORA) on InternLM2 on all $p^{+}$ prompts. Intuitively fine-tuning should learn to produce natural language thought from $p^{+}$. We do LORA with rank $8$, alpha $16$ and learning rate $2e^{-5}$ for 3 epochs. In figure 5 we compare finetuning results with activation steering, we showcase that finetuning InternLM2 yielded worse results than steering vectors. We attribute this behavior due to the limited number of $p^{+}$ prompts. We also made an attempt to fine-tune the model with the original set we sampled from LeanSTaR data.
Cost analysis We conduct all our experiments on Nvidia A100, for LORA finetuning, at rank 8 there are $1,338,656$ parameters. It took around 1 hour for training on our training data. In comparison steering only requires inference time updates making it much more efficient.
Cross Lingual Transfer To understand the true quality of these steering vectors we also evaluate InternLM2 on generating Coq proofs. This task helps us understand if the steering vectors only learn how to generate the natural language information while generating the proof or the core artifacts needed to imporve in Lean. In table 2 we show that steering vectors generate 1 valid proof in Coq from Putnambench. This cross lingual transfer validates the
Figure 5: Comparison between the pass rates achieved from models after doing LORA finetuning with parameters $\{rank=8,\alpha=16,learning\_rate=2\times 10^{-5}\}$ on prompts $p^{+}$. The pass rate are reported at 1x32x100 for InternLM2.
4.2.2 Layer responsibility
To localize where distilled informal thoughts most effectively influence proof generation, we perform layer-wise activation patching on our fine-tuned theorem-proving model. Specifically, we replace the model’s native activation vectors at contiguous layer blocks early (layers 5–13), middle (layers 14–24), and late (layers 25–30)—with those extracted under distilled reasoning prompts, and measure the resulting changes in miniF2F accuracy.
We find that interventions in the early layers yield only modest improvements, suggesting that these layers primarily encode low-level syntactic or symbol-manipulation features. Patching the middle layers leads to intermediate gains, indicating the emergence of conceptual inference patterns. Most notably, injecting distilled activation content into the final transformer blocks produces the largest accuracy boost. This highlights that the deepest layers provide the greatest leverage for steering proof-writing behavior, and that distilled reasoning signals most directly guide final proof synthesis and tactic sequencing.
        Selected Layers Pass Rates
        25-30 51.6%
        14-24 50.8%
        5-13 49.5%
Table 3: Pass rate at $2\times 32\times 600$ for InternLM2.5-StepProver at selected layers on miniF2f. This showcases how the proof construction relies on latter layers of language model and the intial layers are core part to comprehend and represent the theorems in activation space.
4.2.3 Avg Proof Length
To analyze the structural characteristics of the generated proofs, we track the usage of the have tactic, which is commonly used to introduce intermediate lemmas or helper statements. We find that proofs guided by informal reasoning contain a higher frequency of the have tactic, suggesting that such reasoning encourages a modular proof construction with explicitly stated intermediate steps.
Table 5 presents a comparative analysis of the theorems proved by InternLM2.5-StepProver with and without the use of a steering vector. The steering vector notably improves performance on shorter proofs, whereas the gains diminish for longer proofs. We hypothesize that this is due to the informal reasoning becoming less reliable as proof length increases as longer proofs require deeper logical consistency, and informal thoughts may introduce noise or misleading cues, thereby hindering accurate tactic prediction in such cases.
Model Decoding Random Steering
LLEMMA Sampling 22.7% 28.1%
Search 19.2% 26.3%
InternLM2 Sampling 21.4% 32.4%
Search 18.9% 26.8%
InternLM2.5-StepPtover Sampling 40.3% 52.4%
Table 4: Pass rates on miniF2F with randomized vectors and steering vectors.
Proof Length (steps) Without Steering With Steering
<5 76 83
6–10 16 13
10–15 10 11
15–30 15 13
30 < 11 8
Table 5: Proof length analysis on InternLM2.5-StepProver on miniF2F-test
4.2.4 Random Steering Vectors
A potential validity concern with any intervention involving activation patching is that the observed improvements might not stem from genuine performance enhancements but rather from activating backup circuits (McGrath et al.,, 2023; Lucchetti and Guha,, 2024). For instance, patching could merely introduce noise into the embedding space, inadvertently triggering alternative pathways that lead to the desired outcome. This phenomenon complicates the interpretability of both patches and steering vectors. To examine this, we conducted an experiment using a randomly generated steering vector (denoted ”Random” in Table 5). Our findings show that even random steering achieves a nonzero accuracy, albeit significantly lower than that of our computed steering vectors. We hypothesize that this residual accuracy arises due to backup circuits. Nevertheless, the substantially higher performance of our computed steering vectors suggests that our approach induces meaningful transformations toward the correct target.
4.3 How much does steering vectors improve proof writing?
We observed that steering helps to rerank responses which help us reduce the number of samples we expand at each node for Best First Search Tree method. In figure 6 we show how the pass rate improves with varying value of expansion width. This showcases the effectiveness of steering to improve
Figure 6: Pass at varying width (N) with best first search parameters $N\cdot 256\cdot 32$ in miniF2F-Test benchmark, we pick samples that are passed both with and without steering.
5 Conclusion
Our work presents a deep analysis of the underlying mechanisms that drive natural language thoughts in Large Language Models for formal theorem proving. By isolating and manipulating specific activation directions associated with natural language ”thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning. Our experiments showcase how language models comprehend natural language very differently compared to formal proof steps and theorems. We explore the use of activation vectors to represent the task of producing NL information (or thought) with the proof step. And it performs consistently better than base models trained for theorem proving. We provide a closer look into how language models represent theorems in activation space to generate proofs. The use of steering vectors to isolate layers responsible for formal reasoning shows the promise in this approach. We believe this is a strong step towards understanding why language models can not reason on olympiad level problems natively.
6 Limitations and Future Work
This work investigates the use of activation steering to interpret and enhance the performance of language models on theorem-proving tasks. While our explorations are thorough and yield consistent empirical results, there are several limitations worth noting.
First, our study focuses solely on incorporating natural language information to guide the theorem-proving process. However, we do not systematically evaluate the relevance or factual correctness of the natural language inputs with respect to the underlying proof obligations. In certain cases, the model may generate correct tactics even when the provided natural language guidance is partially incorrect or misleading. Understanding this phenomenon requires a deeper analysis of the relationship between instruction quality and proof validity, which we leave for future work.
Appendix A Appendix
In this section, we present a series of representative proof trajectories illustrating both the capabilities and failure modes of language models, and demonstrate how activation steering ameliorates key deficiencies. These examples drawn from Lean and Coq benchmarks reveal not only the quantitative improvements afforded by steering vectors, but also their qualitative impact on proof structure and robustness.
Figure 7: Comparison between problems solved with and without steering.
A.1 MiniF2F Problems Solved
We observe that the incorporation of steering vectors results in the generation of a substantial number of new proofs that are not derivable by the base language model alone. We show a detailed plot for all the problems in minif2f in figure 7.
This improvement in proof discovery is accompanied by a corresponding increase in failure cases. We hypothesize that this phenomenon arises due to biases introduced by the data used to construct the steering vectors. Specifically, the induced vector may encode inductive priors that align well with certain classes of theorems while misaligning with others, leading to erroneous proof trajectories. This trade-off highlights the importance of characterizing and formalizing the inductive biases embedded in steering vectors, particularly when generalizing across diverse proof distributions.
A.2 Examples of proofs
A.2.1 How does Steering help in the proof?
In the example below, we demonstrate how injecting a steering vector introduces the natural language information at each step that helps the LLM to write a proof more robustly. In the example below, we observe the use of high-level overview of the next step which helps in grounding the generation. We also compare it with the unsteered counterpart of the same problem and discuss the key differences in proof writing.
Example: mathd_algebra_156 in miniF2F benchmark.
import Mathlib
import Aesop
set_option maxHeartbeats 0
open BigOperators Real Nat Topology Rat
theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
(h1 : $\forall$ t, g t = 5 * t ^ 2 - 6) (h2 : f x = g x) (h3 : f y = g y) (h4 : x ^ 2 < y ^ 2) :
y ^ 2 - x ^ 2 = 1 := by
-- Apply function definitions to rewrite f x and g x in terms of powers of x
have eq1 : x^4 = 5 * x^2 - 6 := by
rw [h0, h1] at h2
linarith
-- Apply function definitions to rewrite f y and g y in terms of powers of y
have eq2 : y^4 = 5 * y^2 - 6 := by
rw [h0, h1] at h3
linarith
-- Rearranged form: x^4 - 5 x^2 + 6 = 0
-- Use known factorization of the quartic as product of quadratics
have h5 : (x^2 - 2) * (x^2 - 3) = 0 := by
have h1 : x^4 - 5 * x^2 + 6 = 0 := by linarith
have h2 : x^4 - 5 * x^2 + 6 = (x^2 - 2) * (x^2 - 3) := by
ring
nlinarith
-- Apply same reasoning for y using identical polynomial structure
have h6 : (y^2 - 2) * (y^2 - 3) = 0 := by
have h1 : y^4 - 5 * y^2 + 6 = 0 := by linarith
have h2 : y^4 - 5 * y^2 + 6 = (y^2 - 2) * (y^2 - 3) := by
ring
nlinarith
-- Conclude that x^2 must be either 2 or 3 based on zero product property
have hx2 : x^2 = 2 $\vee$ x^2 = 3 := by
cases (mul_eq_zero.mp h5) with
| inl h =>
left
nlinarith
| inr h =>
right
nlinarith
-- Similarly, conclude that y^2 must be either 2 or 3
have hy2 : y^2 = 2 $\vee$ y^2 = 3 := by
cases (mul_eq_zero.mp h6) with
| inl h =>
left
nlinarith
| inr h =>
right
nlinarith
-- Perform case analysis on possible values of x^2 and y^2
cases hx2 with
| inl hx2l =>
-- Case: x^2 = 2
cases hy2 with
| inl hy2l =>
-- Case: y^2 = 2, contradicts h4 : x^2 < y^2
exfalso
nlinarith
| inr hy2r =>
-- Case: y^2 = 3, compute difference
have h : y^2 - x^2 = 1 := by
rw [hy2r, hx2l]
linarith
linarith
| inr hx2r =>
-- Case: x^2 = 3
cases hy2 with
| inl hy2l =>
-- Case: y^2 = 2, contradicts x^2 < y^2
exfalso
nlinarith
| inr hy2r =>
-- Case: y^2 = 3, contradicts x^2 < y^2
exfalso
nlinarith
Comparison with Unsteered Proof The unsteered model demonstrates a breakdown in aligning its reasoning with the algebraic form of the goal. Specifically, it bypasses the canonical application of the zero-product property via mul_eq_zero, which is essential for decomposing polynomial equalities into tractable cases. Instead, it resorts to a general-purpose disjunction strategy using eq_or_ne, which introduces branches without constructive grounding and fails to anchor the analysis in the underlying factorization. This leads to brittle reasoning that lacks semantic traceability to the original assumptions. In contrast, the steered model adheres to a proof structure that mirrors standard mathematical practice: it identifies the factorable structure of the quartic polynomials, applies mul_eq_zero to extract meaningful cases, and maintains the logical continuity required to propagate constraints across steps. The comparison reveals how steering enhances the model’s sensitivity to algebraic invariants and encourages the use of idiomatic tactics, leading to more reliable and interpretable proofs.
Diff of proofs
diff --git a/mathd_algebra_156_internlm2-5_w_steering.lean b/mathd_algebra_156_internlm2-5_wo_steering.lean
index b79f717..7f48d26 100644
--- a/mathd_algebra_156_internlm2-5_w_steering.lean
$+$ b/mathd_algebra_156_internlm2-5_wo_steering.lean
@@ -3 +2,0 @@ import Aesop
@@ -5 +3,0 @@ set_option maxHeartbeats 0
@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t
- have eq1 : x^4 = 5 * x^2 - 6 := by
- rw [h0, h1] at h2
- linarith
- have eq2 : y^4 = 5 * y^2 - 6 := by
- rw [h0, h1] at h3
- linarith
- have h5 : (x^2 - 2) * (x^2 - 3) = 0 := by
- have h1 : x^4 - 5 * x^2 + 6 = 0 := by linarith
- have h2 : x^4 - 5 * x^2 + 6 = (x^2 - 2) * (x^2 - 3) := by
- ring
- nlinarith
- have h6 : (y^2 - 2) * (y^2 - 3) = 0 := by
- have h1 : y^4 - 5 * y^2 + 6 = 0 := by linarith
- have h2 : y^4 - 5 * y^2 + 6 = (y^2 - 2) * (y^2 - 3) := by
- ring
- nlinarith
- have hx2 : x^2 = 2 $\vee$ x^2 = 3 := by
- cases (mul_eq_zero.mp h5) with
- | inl h =>
- left
- nlinarith
- | inr h =>
- right
- nlinarith
- have hy2 : y^2 = 2 $\vee$ y^2 = 3 := by
- cases (mul_eq_zero.mp h6) with
- | inl h =>
- left
- nlinarith
- | inr h =>
- right
- nlinarith
- cases hx2 with
- | inl hx2l =>
- cases hy2 with
- | inl hy2l =>
- exfalso
- nlinarith
- | inr hy2r =>
- have h : y^2 - x^2 = 1 := by
- rw [hy2r, hx2l]
- linarith
- linarith
- | inr hx2r =>
- cases hy2 with
- | inl hy2l =>
- exfalso
- nlinarith
- | inr hy2r =>
+ have hx : x ^ 4 = 5 * x ^ 2 - 6 := by rw [← h0, h1, h2]
+ have hy : y ^ 4 = 5 * y ^ 2 - 6 := by rw [← h0, h1, h3]
+ have eq_x : x ^ 4 - 5 * x ^ 2 + 6 = 0 := by linarith [hx]
+ have eq_y : y ^ 4 - 5 * y ^ 2 + 6 = 0 := by linarith [hy]
+ have factor_x : (x ^ 2 - 2) * (x ^ 2 - 3) = 0 := by
+ calc (x ^ 2 - 2) * (x ^ 2 - 3)
+ = x ^ 4 - 3 * x ^ 2 - 2 * x ^ 2 + 6 := by ring
+ _ = x ^ 4 - 5 * x ^ 2 + 6 := by ring
+ _ = 0 := eq_x
+ have factor_y : (y ^ 2 - 2) * (y ^ 2 - 3) = 0 := by
+ calc (y ^ 2 - 2) * (y ^ 2 - 3)
+ = y ^ 4 - 3 * y ^ 2 - 2 * y ^ 2 + 6 := by ring
+ _ = y ^ 4 - 5 * y ^ 2 + 6 := by ring
+ _ = 0 := eq_y
+ have x_cases : x ^ 2 = 2 $\vee$ x ^ 2 = 3 := by
+ cases’ eq_or_ne (x ^ 2 - 2) 0 with h h
+ · left; linarith [h]
+ · right
+ have : x ^ 2 - 3 = 0 := by
+ have : (x ^ 2 - 2) $\neq$ 0 := h
+ exact eq_of_mul_eq_zero_right factor_x this
+ linarith [this]
+ have y_cases : y ^ 2 = 2 $\vee$ y ^ 2 = 3 := by
+ cases’ eq_or_ne (y ^ 2 - 2) 0 with h h
+ · left; linarith [h]
+ · right
+ have : y ^ 2 - 3 = 0 := by
+ have : (y ^ 2 - 2) $\neq$ 0 := h
+ exact eq_of_mul_eq_zero_right factor_y this
+ linarith [this]
+ cases’ x_cases with hx_2 hx_3
+ · cases’ y_cases with hy_2 hy_3
+ · exfalso
+ rw [hx_2, hy_2] at h4
+ exact lt_irrefl 2 h4
+ · rw [hy_3, hx_2]
+ norm_num
+ · cases’ y_cases with hy_2 hy_3
+ · exfalso
+ rw [hx_3, hy_2] at h4
+ norm_num at h4
@@ -68 +57,2 @@
- nlinarith
+ rw [hx_3, hy_3] at h4
+ exact lt_irrefl 3 h4
A.2.2 Does Steering introduce only useful natural language information?
We analyze the proofs and the natural language incorporation added by steering vectors, and we were surprised to observe that in few cases the natural language instructions were incorrect and the model was able to come up with the correct formal proofs. This shows the capability of choosing the correct step not only depends on the correctness of the thoughts but also at the confidence of the model at the time of generation.
Example of mathd_algebra_313 where the intermediate step is incorrectly annotated.
The failure arises from the premature application of the simplification tactic field_simp immediately after the symbolic rewrite of $i$ as $v/z$, based on the hypothesis $v=i\cdot z$. At this stage of the proof, neither $v$ nor $z$ have been instantiated with their concrete values. Consequently, the simplification is applied to the expression $i=(i\cdot z)/z$, which reduces trivially to $i=i$ under cancellation. While syntactically valid, this reasoning is circular and fails to advance toward the proof goal, which requires computing an explicit value for $i$. Furthermore, such simplification is semantically brittle because it assumes invertibility of $z$ (i.e., $z\neq 0$), which is not yet established or unfolded in the proof context. The appropriate use of field_simp should follow substitution of the concrete definitions for both $v$ and $z$, ensuring the resulting expression is a well-defined complex fraction amenable to symbolic manipulation and numeric normalization.
Lean Proof for mathd_algebra_313
import Mathlib
import Aesop
set_option maxHeartbeats 0
open BigOperators Real Nat Topology Rat
theorem mathd_algebra_313 (v i z : $\mathbb{C}$) (h0 : v = i * z) (h1 : v = 1 + Complex.I
) (h2 : z = 2 - Complex.I) : i = 1 / 5 + 3 / 5 * Complex.I := by
-- rewrite i in terms of v and z, to isolate i from the product
have h3 : i = v / z := by
-- replace v with i x z from hypothesis h0
rw [h0]
-- simplify division by cancelling z terms directly (incorrect)
field_simp [Complex.ext_iff, h2]
-- replace i by the fraction v/z, then rewrite v and z
rw [h3, h1, h2]
-- simplify complex division into real and imaginary parts
field_simp [Complex.ext_iff]
-- normalize numbers to explicit fractions
norm_num
A.3 Transferability of Steering Vectors
To investigate the transferability of steering vectors across formal systems, we evaluated their impact when applied to a Coq-based proof assistant. The primary objective of this evaluation was to assess whether these steering vectors merely encode syntactic preferences for generating natural language interleaved with formal proof steps, or whether they also capture reasoning strategies that generalize across distinct proof environments.
In this setting, we observed that the InternLM2.5 model, when guided by the learned steering vectors, was able to successfully construct a proof for Putnam_1988_B1 in Coq. In contrast, the same model failed to complete the corresponding proof in Lean. This discrepancy suggests that the steering vectors encapsulate transferable reasoning features that are not specific to the prover, possibly due to differences in tactic libraries, proof term verbosity, or the model’s learned prior over formal syntax. The result highlights the non-trivial interaction between the encoded reasoning direction and the target proof environment, underscoring the potential of steering vectors as cross-system inductive biases.
Problem Statement for Putnam_1988_B1
A composite number is a positive integer that can be expressed as a product (ab), where (a) and (b) are (not necessarily distinct) integers from the set ({2, 3, 4, dots}).
Problem: Show that every composite number can be written in the form
xy + xz + yz + 1,
where (x, y, z) are positive integers.
Putnam_1988_B1 proof in Coq
Require Import Arith Lia Ring.
(* A number is composite if it can be written as a product of two integers greater than 1 *)
Definition composite (n: nat) : Prop :=
exists x y, 1 < x /\ 1 < y /\ n = x * y.
Lemma Putnam_1988_B1 n :
3 < n ->
composite n ->
exists x y z : nat,
0 < x /\ 0 < y /\ 0 < z
/\ n = x*y + y*z + z*x + 1.
Proof.
(* Using the definition of composite, obtain integers a and b such that n = a * b with a, b > 1 *)
intros _ [a [b [Ha [Hb Hab]]]].
(* The goal is to express n in the form x*y + y*z + z*x + 1 where x, y, z are positive *)
(* Define x and y as one less than a and b, respectively *)
set (x := a - 1). (* Then a = x + 1 *)
set (y := b - 1). (* Then b = y + 1 *)
(* Set z to 1 in order to satisfy symmetry in the expression *)
set (z := 1).
(* Establish positivity of x, y, z *)
exists x, y, z.
(* Since a > 1, x = a - 1 > 0 *)
split; [ lia | ].
(* Since b > 1, y = b - 1 > 0 *)
split; [ lia | ].
(* z is defined as 1, so clearly positive *)
split; [ lia | ].
(* Now verify the identity n = x*y + y*z + z*x + 1 *)
unfold x, y, z.
(* Replace n by a * b as given in the hypothesis *)
rewrite Hab.
(* Simplify the right-hand side of the target expression *)
simpl.
(* Resolve the final arithmetic equivalence *)
lia.
Qed.

### v5

@@
-Steering LLMs for Formal Theorem Proving
+Steering Language Models for Theorem Proving
 Abstract
-Recent progress in automated theorem proving leverages Large Language Models (LLMs) for their capacity to comprehend informal mathematical statements and generate corresponding formal proofs.
-Even though these techniques perform well, very little exploration has been done to understand how language models interpret and utilize these informal mathematical cues to generate formal proofs more effectively.
-To address this, we explore activation steering, a lightweight, inference-time mechanism that identifies linear directions in a model’s residual activations corresponding to informal “thought” traces, and nudges those activations to improve proof construction entirely without finetuning.
-Unlike previous approaches, activation engineering offers valuable insights into language models’ internal reasoning dynamics encoded in their activation space.
-We evaluated these activation vectors on two distinct tasks: formal proof generation from formal theorems and formal proof generation from informal problem descriptions.
-Our contributions are twofold: (1) we purpose an activation-based intervention technique to guide proof synthesis in LLMs; and (2) improve performance across two different decoding strategies without additional training.
-Machine Learning, ICML
+Recent advances in automated theorem proving use Large Language Models (LLMs) to translate informal mathematical statements into formal proofs.
+However, informal cues are often ambiguous or lack strict logical structure, making it hard for models to interpret them precisely.
+While existing methods achieve strong performance, little is known about how LLMs internally represent informal cues, or how these influence proof generation.
+To address this, we explore activation steering, an inference-time intervention that identifies linear directions in residual activations associated with informal reasoning traces and adjusts them to improve proof construction without fine-tuning.
+This mechanism also yields interpretable information about how reasoning is internally encoded in the activation space of LLMs.
+We test our method for generating formal proofs from already-formalized theorems.
+Our contributions are twofold: (1) a novel activation-based intervention for guiding proof synthesis in LLMs; and (2) demonstration that this intervention improves performance under two decoding strategies (sampling and best-first search) without any further training.
 1 Introduction
-Interactive proof assistants such as Lean (de Moura et al.,, 2015), Isabelle (Wenzel et al.,, 2008), and Coq (Barras et al.,, 1999) enable formal verification of mathematical proofs and software by leveraging specialized programming languages (Avigad,, 2023; Ringer et al.,, 2019).
-Neural theorem proving, which integrates neural models like Large Language Models (LLMs) with these assistants, has emerged as a promising approach to automating formal reasoning and theorem proving (First et al.,, 2023; Polu and Sutskever,, 2020; Polu et al.,, 2022; Yang et al.,, 2023; Welleck,, 2023).
-This integration is intuitive and powerful: proof assistants enforce formal correctness, while neural models predict and suggest logical steps for proof construction.
-A key challenge, however, is generalizing proof writing to scenarios unseen during training.
-Multiple prior studies have demonstrated that informal (natural language, NL) information can enhance theorem proving in language models (Lin et al.,, 2024; Welleck et al.,, 2021, 2022).
-Despite these advances, it remains unclear how natural language information influences a model’s reasoning process and whether it can be systematically leveraged to improve proof-writing capabilities.
-Our work addresses this gap by investigating the mechanistic role of natural language information in LLM reasoning and developing a method to harness it effectively.
-We find that the task of generating natural language mathematical context can be encoded as activation vectors in the model’s activation space.
-Using activation steering (Panickssery et al., 2024b, ; Turner et al.,, 2024; Lucchetti and Guha,, 2024), we capture differences in model behavior driven by natural language inputs, enabling targeted interventions that enhance proof-writing without affecting unrelated aspects of performance.
-In figure 1 we demonstrate how structured natural language interventions improve the quality and the structure of proof.
-In comparison to natural sampling leads to counterfeit and incorrect proofs.
-We evaluate this approach on three open-source LLMs like Lemma (Azerbayev et al.,, 2024), InternLM-2 (Ying et al.,, 2024), and InternLM2.5-StepProver (Wu et al.,, 2024) that are designed for theorem proving through training on mathematics data.
-Our experiments show that steering vectors with targeted NL information consistently improve proof quality across the MiniF2F and PutnamBench benchmarks.
-Our key contributions are:
+Interactive proof assistants such as Lean (de Moura et al., 2015), Isabelle (Wenzel et al., 2008), and Rocq (Barras et al., 1999) provide the infrastructure for formal verification of mathematical proofs and software.
+They require proofs to be expressed in precise formal languages (Avigad, 2023; Ringer et al., 2019).
+Neural theorem proving, combining LLMs with proof assistants, has shown promise in automating reasoning (First et al., 2023; Polu & Sutskever, 2020; Polu et al., 2022; Yang et al., 2023; Welleck, 2023).
+Prior work (Lin et al., 2024; Welleck et al., 2021; 2022) suggests that training model on natural reasoning for proof steps can improve performance, but it remains unclear how such informal language is internally represented or whether it can be reliably leveraged to improve proof generation.
+Building on the observation that proofs often interleave natural-language reasoning with formal steps, we show that informal natural language (NL) context induces distinct activation patterns in LLMs.
+We extract these patterns as steering vectors as in (Panickssery et al., 2024b; Turner et al., 2024; Lucchetti & Guha, 2024), and use them to intervene at inference time.
+These interventions improve proof generation quality, while preserving model behavior in unrelated aspects.
+We test our approach across three theorem-proving oriented LLMs - Lemma (Azerbayev et al., 2024), InternLM-2 (Ying et al., 2024), and InternLM2.5-StepProver (Wu et al., 2024).
+On established benchmarks MiniF2F and PutnamBench, we find that steering via informal language context consistently improves proof quality under multiple decoding strategies.
+Our key contributions are threefold: (1) We provide mechanistic insights into how informal mathematical or natural language context impacts internal reasoning in LLMs for theorem proving.
+(2) We propose a method to extract activation vectors that encode such informal context, and use these vectors to guide proof construction in a structured, grounded way.
+(3) We demonstrate that activation steering yields consistent improvements in proof success rates on MiniF2F and PutnamBench benchmarks under both search- and sampling-based decoding.
+By focusing on steering in activation space, our approach sheds light on the link between informal mathematical language and formal reasoning steps; without needing to fine-tune model weights.
+We base this on the hypothesis that many reasoning features are encoded as (approximately) linear directions in activation space (Mikolov et al., 2013; Elhage et al., 2021; Nanda et al., 2023; Park et al., 2024).
+Although not all features follow perfect linearity (Engels et al., 2025), this assumption has previously enabled methods like concept erasure and steering to work well (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025).
+The rest of this paper is organized as follows.
+In Section 2, we review related work on neural theorem proving, activation steering, and the representation of informal mathematical reasoning in LLMs.
+Section 3 presents our activation steering method: how steering vectors are constructed from informal mathematical contexts, how they are applied at inference time, and how we select layers and intervention strengths.
+Section 5 describes our experimental setup, including models, benchmarks (MiniF2F, PutnamBench), and decoding strategies (sampling vs best-first search).
+Section 6 reports results: we analyze performance gains, the effect of steering vectors on proof success rates, and ablations experiments.
+Finally, in Section 7 we discuss insights into the internal reasoning of the model, limitations of our approach, and potential directions for future work.
+2 Related Work
+This section situates our work in three intersecting strands: automatic theorem proving and auto‐formalization, representation learning in large language models (LLMs), and mechanistic interpretability via activation interventions.
+Automatic Theorem Proving and Auto-formalization Recent progress in automatic theorem proving frames proof search as sequence generation.
+The GPT-f framework (Polu & Sutskever, 2020) trains a language model to map proof states to tactics, combining it with best‐first search to assemble proofs.
+Extensions explore data augmentation via proof transformations or synthesis (Han et al., 2022; Rotella et al., 2025; Wang & Deng, 2020), improved search strategies (Wang et al., 2023), curriculum training (Polu et al., 2022), and retrieval from proof libraries (Yang et al., 2023).
+Systems like LLMStep unify these ideas into usable frameworks (Welleck & Saha, 2023).
+Auto‐formalization complements this by translating informal mathematics (e.g.
+text, sketches) into formal proofs.
+Surveys highlight translation techniques (Wu et al., 2022), while Draft-Sketch-Prove shows LLMs prompted with informal sketches outperform formal-only baselines (Jiang et al., 2023).
+LeanStar (Lin et al., 2024) interleaves informal reasoning with tactic prediction, typically via supervised fine-tuning on synthetic data.
+A central open question remains: How do informal reasoning patterns inform formal proving within a model’s internal representations?
+Our approach explores this via steering vectors in activation space, injecting implicit natural-language “thoughts” at inference to guide proof steps without heavy fine-tuning, probing the latent link between informal intuition and formal reasoning.
+Language Model Representation of Concepts Our method is motivated by prior work demonstrating that task or feature concepts can be represented as linear directions in the activation space of LLMs when given appropriate contextual examples.
+Hendel et al.
+(2023) and Todd et al.
+(2024) show that a context (e.g.
+a prompt) can induce a task embedding in some activation subspace.
+A broader literature examines linear decompositions of features such as truthfulness (Azaria & Mitchell, 2023; Li et al., 2023; Marks & Tegmark, 2024), sentiment (Tigges et al., 2024), harmlessness (Zou et al., 2025; Zheng et al., 2024), sycophancy or alignment steering (Perez et al., 2023; Panickssery et al., 2024a; Sharma et al., 2024), factual knowledge (Gurnee & Tegmark, 2024), and refusal behavior (Arditi et al., 2024).
+Relatedly, unsupervised methods like sparse autoencoders have been used to extract concept directions in hidden space (Bricken et al., 2023; Huben et al., 2024; Templeton et al., 2024).
+These works generally share the hypothesis that LLMs encode high-level features or concepts as (approximately) linear directions in activation space (Elhage et al., 2021; Mikolov et al., 2013; Nanda et al., 2023; Park et al., 2024).
+While recent work cautions that not all features may admit clean linear representations (Engels et al., 2025), the linearity assumption has proven effective in practice for concept erasure, model steering, and interpretability (Beaglehole et al., 2025; Zhao et al., 2025a; Shah et al., 2025).
+Within this framework, we explore whether generating a natural-language informal explanation can itself be represented as a linear direction, and how injecting this direction at inference improves formal theorem proving performance.
+Mechanistic Interpretability and Activation Patching A rich body of work in mechanistic interpretability seeks to locate, analyze, and manipulate internal representations in transformer-based models.
+Early studies localized factual knowledge or associative memory to particular neurons or circuits (Meng et al., 2022), and probed hidden layers for high-level features (Li et al., 2024; Dong et al., 2023).
+The idea of implicit evaluation i.e.
+measuring latent capability of a model rather than just output behavior, has been developed to complement benchmark-based evaluation (Dong et al., 2023).
+One particularly relevant tool is activation patching (or residual stream intervention) (Vig et al., 2020; Variengien & Winsor, 2023), wherein one alters specific activations in a layer (often in the residual stream) at inference time to influence model outputs.
+Recent works have used this to edit factual associations or steer behavior (Zhao et al., 2025b).
+The residual stream is a high-dimensional accumulator of intermediate features (propagated via skip connections) that each transformer layer refines or adds to; intervening there offers a principled way to steer model behavior (Elhage et al., 2021).
+In the domain of theorem proving, activation interventions provide a promising lens into how a model processes and integrates informal guidance with formal reasoning.
+In our approach, we generate steering vectors (patches) that, when applied to the residual stream at certain layers, help the model emit an internal “natural-language thought” prior to or alongside tactic prediction.
+We find that effective steering tends to realign activations such that the output of the model conforms to a structured reasoning format, thereby improving downstream proof search and reducing failure modes.
+Building on these insights, we adapt activation interventions to theorem proving by treating informal reasoning as a direction in activation space.
+We then show how extracting and injecting this direction can guide proof generation.
+The next section details the architecture, design choices, and evaluation of this steering approach.
+3 Steering for improving theorem proving
+We aim to steer theorem-proving models toward using informal reasoning via activation interventions.
+This section describes (i) how we compute steering vectors, (ii) how we choose layers and apply steering, and (iii) efficiency and practical considerations.
+3.1 Intuition and Overview
+Modern transformer models often encode high-level semantic behavior (e.g.
+reasoning, style) as approximately linear directions in activation space (residual streams) (Zou et al., 2025; Elhage et al., 2021).
+We exploit this property: given pairs of prompts that differ only by the presence of natural-language reasoning (but otherwise describe the same proof step), we can estimate a vector in activation space that points in the “informal reasoning” direction.
+At inference time, adding this vector (appropriately scaled) nudges the model toward producing reasoning-augmented proofs.
+In the rest of this section, we detail how we compute these steering vectors, how we pick which layers to intervene on, and how we incorporate them efficiently at inference.
+3.2 Constructing Steering Vectors
+We adopt a difference-of-means (contrastive) method (Belrose et al., 2023), which has been effective at extracting feature directions across multiple domains (e.g.
+refusal, truthfulness) (Arditi et al., 2024; Panickssery et al., 2024b; Marks & Tegmark, 2024).
+Let $\mathcal{D}=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ be a dataset of paired prompts, where $p_{i}$ is a standard formal proof prompt and $p_{i}^{+}$ is its version augmented with explicit natural-language reasoning.
+We feed both prompts into our model $\mathcal{M}$ and extract the residual stream activations at a chosen layer $\ell$.
+Denote the activation at the final token position by $\mathbf{v}_{\ell}^{p}=\text{resid}(\mathcal{M}(p),\ell)$.
+Then the steering vector $\mathbf{u}^{(\ell)}$ is computed as:
+$\mathbf{u}^{(\ell)}=\frac{1}{|\mathcal{D}^{+}|}\sum_{p^{+}\in\mathcal{D}^{+}}\mathbf{v}_{\ell}^{p^{+}}-\frac{1}{|\mathcal{D}^{-}|}\sum_{p^{-}\in\mathcal{D}^{-}}\mathbf{v}_{\ell}^{p^{-}}$ (1)
+Here $\mathcal{D}^{+}$ and $\mathcal{D}^{-}$ are the two halves of the prompt-pair dataset (augmented and unaugmented), so the subtraction isolates the direction most correlated with natural-language reasoning while largely cancelling out shared biases or irrelevant activation patterns.
+Figure 1: Steering Vectors are computed as difference of means for $p^{+}$ and $p$
+We emphasize that this method requires only forward passes; no gradient-based finetuning or parameter updates; making it computationally lightweight.
+3.3 Layer Selection and Intervention Strategy
+Not all layers are equally responsive to steering, so we first perform an activation analysis to find suitable intervention points. For each layer $\ell$, we measure:
+$\text{sim}(\ell)=\frac{1}{|\mathcal{D}|}\sum_{(p,p^{+})\in\mathcal{D}}\cos(\mathbf{v}_{\ell}^{p},\mathbf{v}_{\ell}^{p^{+}})$ (2)
+Empirically, we observe that in early layers, $\text{sim}(\ell)$ is close to 1 (i.e. little divergence), but in deeper layers it drops and exhibits local minima (“valleys”) as seen in Figure 2.
+Figure 2: Cosine similarity across model’s residual stream activations for each layer.
+We find that in early layers, activations remain highly similar between $p$ and $p^{+}$, but as we go deeper, the cosine similarity drops and exhibits local minima.
+Intuitively, these are the layers where the representations diverge most when informal reasoning is introduced (see Figure 2).
+Thus, these “valley” layers are promising candidates for steering.
+These valleys likely correspond to points where the internal representation of the model is most sensitive to reasoning-specific perturbations.
+We select intervention layers where (a) cosine similarity dips significantly, and (b) representation is still semantically stable. At inference, we inject the steering vector into those layers:
+$\mathbf{v}_{\ell}^{\prime}=\mathbf{v}_{\ell}+\alpha\cdot\mathbf{u}^{(\ell)}$ (3)
+We treat $\alpha$ as a hyperparameter and specify that it will be tuned on held-out validation data.
+The method instantiates a sweep to find the optimal balance between influencing reasoning and maintaining validity.
+We place constraints (e.g.
+an upper bound on $\alpha$) to avoid destabilizing proofs.
+We present the complete pseudocode of our approach in Algorithm 1.
+Algorithm 1 Activation Steering for Theorem Proving
+0: Contrastive prompt pairs $D=\{(p_{i},p_{i}^{+})\}_{i=1}^{N}$ where $p_{i}$ are standard prompts and $p_{i}^{+}$ include natural language reasoning
+0: Language model $M$ with $L$ layers
+0: Test theorem $T$ to prove
+0: Enhanced proof generation
+1:Offline: Create Steering Vectors
+2:for each layer $\ell\in\{1,\ldots,L\}$ do
+3:  $\bar{v}^{+}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}^{+}),\ell)$ {Mean activation for augmented prompts}
+4:  $\bar{v}^{-}\leftarrow\frac{1}{N}\sum_{i=1}^{N}\text{resid}(M(p_{i}),\ell)$ {Mean activation for standard prompts}
+5:  $u^{(\ell)}\leftarrow\bar{v}^{+}-\bar{v}^{-}$ {Steering vector at layer $\ell$}
+6:end for
+[…]
+8:Online: Apply Steering During Inference
+9: Initialize proof context with theorem $T$
+10:while proof incomplete and within computational budget do
+11:  Perform forward pass through model
+12:  for each intervention layer $\ell\in\mathcal{L}$ do
+13:   $v^{\prime}_{\ell}\leftarrow v_{\ell}+\alpha\cdot u^{(\ell)}$ {Inject steering vector}
+14:  end for
+15:  Generate next proof step using modified activations
+16:  Validate and apply proof step if correct
+17:end while
+18:return Generated proof or failure
+4 Experimental Setup
+4.1 Models
+We conduct experiments on three open-source 7B-parameter language models optimized for mathematical reasoning:
+  • Llemma-7B (Azerbayev et al., 2024): Built on Code Llama architecture, pre-trained on mathematical texts and formal mathematics corpora.
+  • InternLM2-7B (Ying et al., 2024): LLaMA-based architecture with extended training on mathematical problem-solving.
+  • InternLM2.5-StepProver (Wu et al., 2024): Enhanced variant with expert iteration on large-scale Lean problems.
+All models share transformer architectures with consistent residual stream dimensionality, facilitating uniform steering vector extraction.
+4.2 Data and Vector Construction
+We construct steering vectors from the Lean-STaR dataset(Lin et al., 2024), randomly sampling 10,000 theorem-proof pairs.
+Each datapoint $(p,p^{+})$ consists of a formal proof step $p$ and its augmented version $p^{+}$ containing explicit natural language reasoning.
+For robustness, we generate model responses $r_{p}$ and $r_{p^{+}}$ for each prompt pair and retain only instances where both responses are valid proof steps with $r_{p}\neq r_{p^{+}}$.
+This filtering yields approximately 7,400 high-quality contrastive pairs.
+Vector construction requires only forward passes through the model, consuming $\sim 1$ GPU-hour on NVIDIA A100, a 60× reduction compared to LoRA fine-tuning while achieving superior performance gains.
+4.3 Chosen Layers and $\alpha$
+Based on activation similarity analysis (Section 3.3), we apply steering vectors to Layers 22, 25 in InternLM2-7B and Layers 24, 25 in Llemma-7B.
+These correspond to layers where the model’s representation diverges most from baseline while maintaining semantic coherence.
+We empirically determine $\alpha=0.8$ through validation experiments.
+4.4 Evaluation Benchmarks
+miniF2F Our primary evaluation utilizes miniF2F (Zheng et al., 2022), a standardized benchmark comprising 244 theorems drawn from mathematical competitions (AMC, AIME, IMO).
+These problems span algebra, number theory, geometry, and combinatorics with varying proof complexities.
+We employ two proof search strategies:
+  • Best-first search: Expansion budget $N\in\{50,600\}$, tactic sampling width $S=32$, selecting states by cumulative log-probability
+  • Parallel sampling: $K$ independent proof attempts to accommodate probability shifts from natural language injection
+PutnamBench We additionally evaluate on PutnamBench (Tsoukalas et al., 2024), containing problems from the William Lowell Putnam Mathematical Competition.
+Following standard evaluation protocol, we test on both the Lean subset (657 problems) and Rocq subset (412 problems) using the benchmark’s default search parameters: $N=600$ expansion budget with $S=32$ tactic width for best-first search, and parallel sampling with $K=2$ attempts.
+This benchmark features more challenging problems with longer average proof lengths, providing a stringent test of steering effectiveness on complex mathematical reasoning.
+Evaluation Metrics.
+Following established practice in neural theorem proving, we report pass rates (percentage of theorems successfully proved within computational budget) as our primary metric.
+For detailed analysis, we additionally examine proof characteristics including average length, tactic distribution, and the frequency of intermediate lemma usage (via the have tactic) to understand how steering affects proof structure and strategy.
+5 Results and Analysis
+Our experiments were designed to address the following research questions:
   1. 1.
-    Mechanistic insights into how informal mathematical context and natural language information affects LLM reasoning about formal theorems.
+    Does activation steering improve theorem-proving performance of existing models?
   2. 2.
-    Activation vector identification for encoding natural language information to guide formal proof construction in a structured and grounded manner.
+    Which layers contribute most effectively when steered?
   3. 3.
-    Activation steering framework yielding consistent gains in proof generation quality on MiniF2F and PutnamBench.
-This study not only improves how models reason about specific tasks, but also provides a practical, interpretable tool for steering reasoning.
-These learning’s provide crucial insights to interpret and understand LLMs for theorem proving better.
-2 Related Work
-Automatic Theorem Proving and Auto-formalization Prior work on automatic theorem proving typically follows the GPT-f framework (Polu and Sutskever,, 2020), which trains a language model on proof state and next tactic pairs, then proves theorems using this model with best-first tree search.
-Subsequent work has explored several directions, including data augmentation (Han et al.,, 2022), synthetic data for training language models (Rotella et al.,, 2025; Wang and Deng,, 2020), novel proof search methods (Wang et al.,, 2023), further training through curriculum learning (Polu et al.,, 2022), retrieval augmented generation (Yang et al.,, 2023), and practical tools (Welleck and Saha,, 2023).
-Figure 1: In this example we show how steering vectors introduce natural language grounding in the proof improving the quality.
-We have highlighted the effect of steering vectors in yellow.
-The lines highlighted cyan represent incorrect tactics leading to incorrect proof.
-A parallel thread focusing on improving automated theorem proving has explored auto-formalization (Wu et al.,, 2022), which incorporates informal mathematics into formal theorem proving.
-This direction recognizes the significant value of connecting informal mathematical intuition with formal proof systems.
-Draft-Sketch-Prove (Jiang et al.,, 2023) demonstrates that language models can leverage informal proofs to enhance formal proving abilities through few-shot prompting rather than supervised fine-tuning.
-Similarly, LeanStar (Lin et al.,, 2024) introduces ”thought-augmented reasoning” in each proof step, using informal reasoning to guide tactic prediction.
-However, approaches like LeanStar typically require supervised fine-tuning of large-language models on synthetic data to improve reasoning capabilities.
-Our work addresses the fundamental challenge of understanding the relationship between informal and formal mathematics by learning task of producing thoughts directly in activation space to improve theorem proving at inference time.
-This approach provides a lightweight alternative to learning informal reasoning patterns through fine-tuning and offers insights into how language models internally represent and connect informal mathematical reasoning with formal proof steps.
-Understanding this relationship is crucial for developing more effective theorem proving systems that can leverage the complementary strengths of both informal mathematical intuition and formal verification.
-Language model representation Our approach is inspired by previous research that shows that it is possible to obtain vector representation of tasks that language model learns in context (Hendel et al.,, 2023; Todd et al.,, 2024).
-These studies are part of broader body of work that examines the linear representation of features such as truthfulness (Azaria and Mitchell,, 2023; Li et al.,, 2023; Marks and Tegmark,, 2024), sentiment (Tigges et al.,, 2024), harmlessness (Zou et al.,, 2023; Zheng et al.,, 2024), sychophancy (Perez et al.,, 2023; Panickssery et al., 2024a, ; Sharma et al.,, 2024), factual knowledge (Gurnee and Tegmark,, 2024), and refusal (Arditi et al.,, 2024).
-In addition, recent works have employed sparse autoencoders to identify feature directions in an unsupervised manner (Bricken et al.,, 2023; Huben et al.,, 2024; Templeton et al.,, 2024).
-A shared hypothesis across these works is that LLMs represent features or concepts as linear directions in activation space (Mikolov et al.,, 2013; Elhage et al.,, 2021; Nanda et al.,, 2023; Park et al.,, 2024).
-While recent studies suggest that not all features may be linearly encoded (Engels et al.,, 2025), the linearity assumption has been effective for both concept erasure and model steering.
-In our work we explore how the concept of generating natural language informal data is a linearly representable concept in activation space which improves the performance of these models to prove formal theorems.
-Mechanistic Interpretability and Activation Patching Previous work has focused on localizing and editing factual associations within transformers (Meng et al.,, 2022) and probing hidden representations for high-level knowledge (Li et al.,, 2024; Dong et al.,, 2023).
-Such studies show implicit evaluations the ability of the model, complementing explicit benchmarks (Dong et al.,, 2023).
-A key technique in mechanistic interpretability is activation patching (Vig et al.,, 2020; Variengien and Winsor,, 2023), which modifies model activations in the residual stream at the time of inference to influence outputs (Zhao et al.,, 2025).
-Residual streams is the high-dimensional vector that carries and accumulates information across layers via skip (residual) connections allowing each layer to iteratively refine and add shared representation used for prediction.
-In figure 2 we demonstrate the process of generating activation vectors on the proof step generation task to enable language models to generate natural language patch that contains context for the next tactic.
-For theorem proving, mechanistic interpretability should provide insight onto how well language models can comprehend natural language information and use it to comprehend formal proofs.
-By dissecting these representations, we can identify failure cases, refine step generation, and enhance the proof generation capability.
-We observe that effective steering transforms activations to align the model’s responses to a structured format to generate natural language information while writing a proof step that allows them to be better at proofing theorems.
-3 Methodology
-3.1 Model Choice
-We build steering datasets for 7B parameter Llemma (Azerbayev et al.,, 2024), InternLM2 (Ying et al.,, 2024) and InternLM2.5-StepProver (Wu et al.,, 2024).
-These models are trained to generate formal theorems and on popular programming languages to improve their capability of reason about theorem’s better.
-InternLM family of models are based on LlaMA architecture (Touvron et al.,, 2023), however Llemma model is trained on CodeLlaMA architecture.
-Figure 2: Steering Vectors are computed as difference of means for for $p^{+}$ and $p$
-3.2 Tasks and Source Data
-We constructed steering pairs from a randomly sampled subset $S$ of the Lean-STaR dataset  [^1] https://huggingface.co/datasets/ScalableMath/Lean-STaR-plus (Lin et al.,, 2024).
-This subset consists of $10,000$ unique data points.
-Each of these data points consist of a formal proof step and an informal thought required to generate that step.
-Every natural prompt $p\in S$ is a formal proof step of a proof.
-We then create $p^{+}$ by inserting explicit reasoning step of natural language explanations of the proof state.
-This process gives us prompts pairs that are required to generate steering vectors.
-After generating prompt pairs $(p,p^{+})$, we prompt the model to with both the prompts and collect responses $r_{p}$ and $r_{p^{+}}$.
-We then filter a subset $s\subset S$ where both $r_{p}$ and $r_{p^{+}}$ are valid proof steps and $r_{p}\neq r_{p^{+}}$.
-The resulting quadruples $\{p,p^{+},r_{p},r_{p^{+}}\}$ form the basis for steering vector construction.
-Formal Theorem to Formal Proof Writing In this task, the model has generate a proof step-by-step for a given formal problem.
-We use standard benchmark for this task miniF2F (Zheng et al.,, 2022).
-This benchmark contains 244 test set theorems in Lean.
-To generate proof for the problem we use best first search as well as beam search.
-We describe the experiment setup in Section 4.
-Informal Problem to Formal Proof Writing This task covers writing proof from informal problem statement.
-We use PUTNAMBENCH (Tsoukalas et al.,, 2024) as the benchmark for this task.
-This benchmark includes 657 problems in Lean and 640 problems in Coq.
-We use both these datasets to evaluate steering vectors.
-3.3 Constructing Steering Vectors
-To extract robust steering vectors, we first identify the activations where these vectors are linearly represented within the model.
-We focus on residual stream activations i.e the outputs of each transformer layer.
-Given an informal thought step we identify the layers where the activation vector is active at time of processing the last token.
-To identify the causally relevant layers, we first extract steering vector candidate from every layer using Difference of Means method.
-This is a widely used technique (Turner et al.,, 2024; Arditi et al.,, 2024).
-This technique is based on constructing contrastive dataset that differs in a specific concept.
-For our task the contrastive dataset is $s$ and the concept we want to differentiate is producing NL information with the formal proof-step.
-We begin with a data set of pairs of prompts and their responses $\{(p_{i},p_{i}^{+},r_{p_{i}},r_{p_{i}^{+}})\}_{i=1}^{N}$, where $p_{i}^{+}$ includes informal reasoning steps and $p_{i}$ is the standard prompt.
-The model $M$ and extract its residual stream activation at the final token of layer $\ell$, denoted
-$v_{p_{i}^{-}}^{\ell}=\mathrm{resid}(M(p_{i}^{-}),\ell)\quad\text{and}\quad v_{p_{i}^{+}}^{\ell}=\mathrm{resid}(M(p_{i}^{+}),\ell).$
-To isolate the component capturing natural language information about the proof, we apply the Difference of Means method:
-$u^{(\ell)}=\frac{1}{|s_{+}|}\sum_{p^{+}\in S_{+}}v_{p^{+}}^{\ell}\;-\;\frac{1}{|s_{-}|}\sum_{p^{-}\in S_{-}}v_{p^{-}}^{\ell},$
-Where $v_{p}^{\ell}$ represents the activation vector of layer $\ell$ for a prompt $p$ and $s_{+}$ represents all the samples in the set $s$ that exhibit the targeted behavior while $s_{-}$ represent all the samples.
-In this scenario the Difference of Means vector is computed by subtracting overall mean activation from the mean activation of the behavior associated with the example.
-This isolates the direction in the activation space most associated with the target behavior while reducing the influence of general model biases.
-Figure 3: Cosine similarity in InternLM2’s residual stream activations drops in later layers for informal reasoning queries, but gradually recovers.
-4 Results and Analysis
-Setup We construct proofs using best-first search method which is one of the most popular methods to evaluate the theorem proving ability of a language model (Polu and Sutskever,, 2020; Yang et al.,, 2023; Azerbayev et al.,, 2023; Lin et al.,, 2024).
-For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$.
-Following standard practice (Polu and Sutskever,, 2020; Yang et al.,, 2023; Welleck and Saha,, 2023; Lin et al.,, 2024), we assume that the state with maximum negative log-probabilities is the ”best” state.
-Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average logarithmic probability of each generated token.
-We expand up to $N$ states and we get a successful proofs search when we reach any proof state with no goals.
-We observed that the best first search perform poorly in our case as with the natural language tokens in our response the mean of log probabilities is noisy and inconsistent, making it harder to choose the post promising branch.
-This observation is consistent to Lean-STaR (Lin et al.,, 2024).
-To overcome these limitation we apply Sampling technique proposed by Lean-STaR (Moura and Ullrich,, 2021).
-Sampling is based on sampling trajectories in parallel for each problem, each time generating at most N tactics.
-The illegal sampled tactics are ignored during sampling.
-This means at state $s_{i}$ if the sampled tactic is $a_{i}$ is only added to trajectory $(s_{0},a_{0}),(s_{1},a_{1}),...,(s_{i-1},a_{i-1})$ if $a_{i}$ is a legal tactic otherwise we sample another tactic from the model M.
-This technique takes approximately same time on GPU as best-first search but accommodates natural language thoughts.
-Figure 4: Effect of patched layers on proof generation in miniF2F-test using InternLM2. Y-axis: verified proofs; X-axis: number of patched layers.
-Model Decoding N K S MiniF2F
-LLEMMA-7B SEARCH 50 1 32 26.2%
-InternLM2-7B SEARCH 50 1 32 30.3%
-InternLM2-7B (SFT) SEARCH 50 1 32 30.7%
-LEAN-COT (InternLM2-7B) SAMPLING 50 32 1 27.0%
-LEAN-COT (InternLM2-7B) SEARCH 50 1 32 25.4%
-LEAN-STAR (InternLM2-7B) SAMPLING 50 32 1 29.1%
-LEAN-STAR (InternLM2-7B) SEARCH 50 1 32 26.2%
-COPRA (WITH GPT-4) CUSTOMIZED - 60 1 29.9%
-InternLM-2-Math-Plus SEARCH 100 1 32 43.4%
-InternLM-2-StepProver SEARCH 100 1 32 48.8%
-InternLM-2.5-StepProver SEARCH 600 2 32 50.7%
-DeepSeekProver-V1.5 SEARCH 400 32 16 63.5%
-InternLM-2.5-StepProver SEARCH 600 256 32 65.9%
-LLEMMA-7B + Activation vector SAMPLING 50 32 1 28.1%
-LLEMMA-7B + Activation vector SEARCH 50 1 32 26.3%
-InternLM2-7B + Activation vector SAMPLING 50 32 1 32.4%
-InternLM2-7B + Activation vector SEARCH 50 1 32 26.8%
-InternLM-2.5-StepProver + Activation Vector SEARCH 600 2 32 49.18%
-InternLM-2.5-StepProver + Activation Vector SAMPLING 600 2 32 52.4%
-InternLM-2.5-StepProver + Activation Vector SEARCH 600 256 32 65.57%
-InternLM-2.5-StepProver + Activation Vector SAMPLING 600 256 32 66.39%
-Table 1: Pass rates on the miniF2F-test dataset with Lean.
-S is the number of tactics attempted at each expanded node (assumed to be 1 in sampling), and K is the total number of search or sampling attempts per problem.
-Model Sampling Pass at
-Problems Solved
-Lean4 (Out of 657)
-InternLM2 7B SEARCH 4096 4
-InternLM2.5-StepProver SEARCH 2x32x600 6
-Goedel-Prover-SFT SEARCH 512 7
-ABEL SEARCH 596 7
-Self-play Theorem Prover SEARCH 3200 8
-Kimina-Prover-7B-Distill SEARCH 192 10
-InternLM2 7B + Activation Steering SAMPLING 4096 4
-InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 7
-Coq (out of 412)
-InternLM2.5-StepProver SAMPLING 2x32x600 0
-InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 1
-Table 2: Results on PutnamBench with Lean and Coq.
-4.1 Main results
-We construct distinct steering vectors for each model variant, enabling us to isolate and analyze the activation-level effects of distilled informal reasoning.
-A detailed comparison of activation patterns across models is presented in Appendix.
-miniF2F Our results on the miniF2F test set are summarized in Table 1.
-Steering the model’s activations consistently improves performance across decoding strategies when compared to the base model.
-Notably, we observe substantial gains in pass rates under both best-first search and sampling strategies.
-One exception is a slight drop in performance for InternLM-2.5-StepProver under best-first search with activation steering.
-We hypothesize that this is due to the injected natural language content lowering the mean log-probability of tokens, which can affect deterministic decoding strategies more than stochastic ones.
-PutnamBench Table 2 reports results on the Lean and Coq subsets of PutnamBench.
-Our findings further support the effectiveness of activation steering: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Coq problem—an improvement over the base model’s zero success rate.
-This cross-system transfer underscores the generality and robustness of the distilled reasoning signal encoded in the steering vectors.
-4.2 Research Questions
-4.2.1 How effective and efficient are activation steering vectors?
-Finetuning To understand the quality of our steering vectors, we do Low Rank Adapter fine-tuning (LORA) on InternLM2 on all $p^{+}$ prompts.
-Intuitively fine-tuning should learn to produce natural language thought from $p^{+}$.
-We do LORA with rank $8$, alpha $16$ and learning rate $2e^{-5}$ for 3 epochs.
-In figure 5 we compare finetuning results with activation steering, we showcase that finetuning InternLM2 yielded worse results than steering vectors.
-We attribute this behavior due to the limited number of $p^{+}$ prompts.
-We also made an attempt to fine-tune the model with the original set we sampled from LeanSTaR data.
-Cost analysis We conduct all our experiments on Nvidia A100, for LORA finetuning, at rank 8 there are $1,338,656$ parameters.
-It took around 1 hour for training on our training data.
-In comparison steering only requires inference time updates making it much more efficient.
-Cross Lingual Transfer To understand the true quality of these steering vectors we also evaluate InternLM2 on generating Coq proofs.
-This task helps us understand if the steering vectors only learn how to generate the natural language information while generating the proof or the core artifacts needed to imporve in Lean.
-In table 2 we show that steering vectors generate 1 valid proof in Coq from Putnambench.
-This cross lingual transfer validates the
-Figure 5: Comparison between the pass rates achieved from models after doing LORA finetuning with parameters $\{rank=8,\alpha=16,learning\_rate=2\times 10^{-5}\}$ on prompts $p^{+}$.
-The pass rate are reported at 1x32x100 for InternLM2.
-4.2.2 Layer responsibility
-To localize where distilled informal thoughts most effectively influence proof generation, we perform layer-wise activation patching on our fine-tuned theorem-proving model.
-Specifically, we replace the model’s native activation vectors at contiguous layer blocks early (layers 5–13), middle (layers 14–24), and late (layers 25–30)—with those extracted under distilled reasoning prompts, and measure the resulting changes in miniF2F accuracy.
-We find that interventions in the early layers yield only modest improvements, suggesting that these layers primarily encode low-level syntactic or symbol-manipulation features.
-Patching the middle layers leads to intermediate gains, indicating the emergence of conceptual inference patterns.
-Most notably, injecting distilled activation content into the final transformer blocks produces the largest accuracy boost.
-This highlights that the deepest layers provide the greatest leverage for steering proof-writing behavior, and that distilled reasoning signals most directly guide final proof synthesis and tactic sequencing.
-        Selected Layers Pass Rates
-        25-30 51.6%
-        14-24 50.8%
-        5-13 49.5%
-Table 3: Pass rate at $2\times 32\times 600$ for InternLM2.5-StepProver at selected layers on miniF2f.
-This showcases how the proof construction relies on latter layers of language model and the intial layers are core part to comprehend and represent the theorems in activation space.
-4.2.3 Avg Proof Length
-To analyze the structural characteristics of the generated proofs, we track the usage of the have tactic, which is commonly used to introduce intermediate lemmas or helper statements.
-We find that proofs guided by informal reasoning contain a higher frequency of the have tactic, suggesting that such reasoning encourages a modular proof construction with explicitly stated intermediate steps.
-Table 5 presents a comparative analysis of the theorems proved by InternLM2.5-StepProver with and without the use of a steering vector.
-The steering vector notably improves performance on shorter proofs, whereas the gains diminish for longer proofs.
-We hypothesize that this is due to the informal reasoning becoming less reliable as proof length increases as longer proofs require deeper logical consistency, and informal thoughts may introduce noise or misleading cues, thereby hindering accurate tactic prediction in such cases.
-Model Decoding Random Steering
-LLEMMA Sampling 22.7% 28.1%
-Search 19.2% 26.3%
-InternLM2 Sampling 21.4% 32.4%
-Search 18.9% 26.8%
-InternLM2.5-StepPtover Sampling 40.3% 52.4%
-Table 4: Pass rates on miniF2F with randomized vectors and steering vectors.
-Proof Length (steps) Without Steering With Steering
-<5 76 83
-6–10 16 13
-10–15 10 11
-15–30 15 13
-30 < 11 8
-Table 5: Proof length analysis on InternLM2.5-StepProver on miniF2F-test
-4.2.4 Random Steering Vectors
-A potential validity concern with any intervention involving activation patching is that the observed improvements might not stem from genuine performance enhancements but rather from activating backup circuits (McGrath et al.,, 2023; Lucchetti and Guha,, 2024).
-For instance, patching could merely introduce noise into the embedding space, inadvertently triggering alternative pathways that lead to the desired outcome.
-This phenomenon complicates the interpretability of both patches and steering vectors.
-To examine this, we conducted an experiment using a randomly generated steering vector (denoted ”Random” in Table 5).
-Our findings show that even random steering achieves a nonzero accuracy, albeit significantly lower than that of our computed steering vectors.
-We hypothesize that this residual accuracy arises due to backup circuits.
-Nevertheless, the substantially higher performance of our computed steering vectors suggests that our approach induces meaningful transformations toward the correct target.
-4.3 How much does steering vectors improve proof writing?
-We observed that steering helps to rerank responses which help us reduce the number of samples we expand at each node for Best First Search Tree method.
-In figure 6 we show how the pass rate improves with varying value of expansion width.
-This showcases the effectiveness of steering to improve
-Figure 6: Pass at varying width (N) with best first search parameters $N\cdot 256\cdot 32$ in miniF2F-Test benchmark, we pick samples that are passed both with and without steering.
+    Does steering enhance proof structure and search efficiency?
+  4. 4.
+    Are improvements consistent across different search budgets?
+  5. 5.
+    How effective and efficient are steering vectors compared to LoRA fine-tuning?
+  6. 6.
+    Do steering vectors generalize across provers?
+5.1 Q1: Impact of Activation Steering
+Activation steering consistently improves model performance across all tested configurations, enabling the discovery of proofs not derivable by the base models alone.
+Model MiniF2F
+Llemma 7B 26.6%
+InternLM2-7B 28.7%
+InternLM2.5-Step Prover 48.2%
+LLEMMA-7B + Steering 28.1%
+InternLM2-7B + Steering 32.4%
+InternLM2.5-StepProver + Steering 66.4%
+Table 1: Performance on MiniF2F (sampling decoding, 50×32×1).
+Model PutnamBench
+InternLM2 7B 4 (0.6%)
+InternLM2.5-StepProver 6 (0.9%)
+InternLM2 7B + Steering 4 (0.6%)
+InternLM2.5-StepProver + Steering 7 (1.1%)
+Table 2: Performance on PutnamBench (Lean). Results reported out of 657 attempts.
+MiniF2F.
+As shown in Table 2, steering improves performance across all models on the miniF2F-Test benchmark (Zheng et al., 2022).
+We adopt the sampling decoding strategy from LeanStar (Lin et al., 2024), which mitigates variance from natural-language generation and provides more effective node selection than standard Best-First Tree Search (see Appendix for Best-First results).
+Steering introduces structured natural-language comments that enhance mathematical reasoning and proof generation.
+Importantly, we find that steering enables a significant number of new proofs beyond the base model’s reach, suggesting that steering vectors guide the model toward otherwise inaccessible solution paths.
+Interestingly, models occasionally succeed despite incorrect informal reasoning, implying that steering exerts influence beyond surface-level natural language.
+While steering increases successful proofs, it also introduces more failure cases.
+We hypothesize this stems from biases in the data used to construct steering vectors, which encode inductive priors that align well with some theorem classes but misalign with others.
+Despite this trade-off, the net effect is strongly positive, establishing activation steering as a lightweight, parameter-free alternative to fine-tuning.
+PutnamBench.
+On PutnamBench (Table 2), steering improves success rates for both InternLM2 and InternLM2.5 on Lean, and achieves non-trivial gains on Rocq problems (details in Section A.3).
+These results highlight steering’s ability to support reasoning in more complex, multi-step proofs.
+5.2 Q2: Layer Choice
+We evaluate layer sensitivity following the strategy described in Section 3.3.
+Steering vectors are patched layer-wise to analyze their influence.
+Figure 4 and Table 4 show that later layers exert stronger influence on theorem-proving performance.
+Selected Layers Pass Rate (%)
+25–30 (Late) 51.6
+14–24 (Middle) 50.8
+5–13 (Early) 49.5
+Figure 3: Pass rates at $2\times 32\times 600$ for InternLM2.5-StepProver on miniF2F.
+Figure 4: Layer-wise ablation results.
+These findings suggest that steering vectors primarily affect later-stage reasoning circuits.
+To ensure robustness, we also compare against random steering vectors, confirming that observed improvements arise from semantically meaningful directions rather than incidental noise.
+5.3 Q3: Proof Structure and Search Efficiency
+Table 3 shows that steering primarily benefits shorter proofs (<<5 steps), while gains diminish for longer ones (>>15 steps).
+Qualitative analysis suggests that informal reasoning introduces noisy intermediate steps for long proofs, reducing effectiveness.
+Interestingly, steering sometimes shortens proof length by guiding the model toward alternate proof strategies.
+Proof Length Without Steering With Steering (total)
+<<5 76 83 (94)
+6–10 16 13 (19)
+10–15 10 11 (20)
+15–30 15 13 (18)
+>>30 11 8 (11)
+Table 3: Proof length analysis for InternLM2.5-StepProver on miniF2F.
+While new proofs increase substantially, so do failure cases, likely reflecting inductive biases encoded in steering vectors.
+Characterizing these biases will be crucial for extending steering to broader proof distributions.
+5.4 Q4: Search Budget
+Figure 5 shows that steering yields consistent improvements across all search budgets $N$.
+Gains grow with larger budgets: at $N=100$, steering triples performance (6 vs.
+2 proofs), and at $N=600$, steering delivers a 38% improvement (162 vs.
+117 proofs).
+These results suggest that steering vectors not only boost baseline performance but also scale effectively with additional computational resources by guiding the search toward more productive proof trajectories.
+Figure 5: Pass rates on miniF2F with varying search budgets.
+5.5 Q5: Effectiveness vs. LoRA Fine-Tuning
+We compare steering against LoRA fine-tuning (Hu et al., 2021) on InternLM2.
+LoRA models were trained on $p^{+}$ prompts with (rank=8, $\alpha_{\text{LoRA}}$=16) and (rank=32, $\alpha_{\text{LoRA}}$=64).
+Performance is reported on miniF2F after each epoch (Figure 6).
+LoRA with higher rank outperforms steering, but lower-rank adaptations underperform, highlighting their limited representational capacity.
+In contrast, steering achieves competitive performance immediately, without additional training or parameters.
+This demonstrates that activation steering offers a highly parameter-efficient alternative, complementing but not replacing fine-tuning.
+Figure 6: Comparison of steering and LoRA fine-tuning.
+5.6 Q6: Generalization Across Provers
+We evaluate if the steering vector is generalizable across different provers.
+In particular we use Putnambench (Tsoukalas et al., 2024) for evaluating cross lingual transfer for steering vectors.
+A particularly compelling finding emerges from our cross-system evaluation: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Rocq problem an improvement over the base model’s zero success rate.
+Specifically, InternLM2.5 with steering successfully constructs a proof for Putnam 1988 B1 in Rocq, while the same model fails to complete the corresponding proof in Lean.
+In Appendix A.3 we showcase the proof written by InternLM2.5-StepProver.
+This preliminary cross-system success suggests that steering vectors may capture reasoning patterns transferable across systems.
+The transferability suggests that our approach enables model to capture deeper mathematical concepts that transcend the specific formal language, encoding universal reasoning strategies that benefit theorem proving across different proof assistants.
+This finding has important implications for developing general-purpose mathematical reasoning systems, indicating that steering vectors trained on one formal system may provide value across the broader ecosystem of interactive theorem provers.
 6 Conclusion
 Our work presents a deep analysis of the underlying mechanisms that drive natural language thoughts in Large Language Models for formal theorem proving.
-By isolating and manipulating specific activation directions associated with natural language ”thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning.
+By isolating and manipulating specific activation directions associated with natural language “thoughts”, we have demonstrated a method to enhance LLMs’ capabilities in mathematical theorem proving without requiring costly fine-tuning.
 Our experiments showcase how language models comprehend natural language very differently compared to formal proof steps and theorems.
@@
 The use of steering vectors to isolate layers responsible for formal reasoning shows the promise in this approach.
-We believe this is a strong step towards understanding why language models can not reason on olympiad level problems natively.
+We believe this provides insight into the challenges LLMs face on Olympiad-level problems and how activation steering may mitigate them.
 7 Limitations and Future Work
@@
 These examples drawn from Lean and Coq benchmarks reveal not only the quantitative improvements afforded by steering vectors, but also their qualitative impact on proof structure and robustness.
-Figure 7: Comparison between problems solved with and without steering.
 A.1 MiniF2F Problems Solved
@@
 We show a detailed plot for all the problems in miniF2F in Figure 7.
+Figure 7: Comparison between problems solved with and without steering.
 This improvement in proof discovery is accompanied by a corresponding increase in failure cases.
@@
 We also compare it with the unsteered counterpart of the same problem and discuss the key differences in proof writing.
-Example: mathd_algebra_156 in miniF2F benchmark.
+Example: Mathd Algebra 156 in miniF2F benchmark.
 import Mathlib
@@
 open BigOperators Real Nat Topology Rat
-theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
+theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t ^ 4)
 (h1 : $\forall$ t, g t = 5 * t ^ 2 - 6) (h2 : f x = g x) (h3 : f y = g y) (h4 : x ^ 2 < y ^ 2) :
@@
 --- a/mathd_algebra_156_internlm2-5_w_steering.lean
-$+$ b/mathd_algebra_156_internlm2-5_wo_steering.lean
++++ b/mathd_algebra_156_internlm2-5_wo_steering.lean
 @@ -3 +2,0 @@ import Aesop
 @@ -5 +3,0 @@ set_option maxHeartbeats 0
-@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : $\forall$ t, f t = t
+@@ -11,56 +9,47 @@ theorem mathd_algebra_156 (x y : $\mathbb{R}$) (f g : $\mathbb{R}$ $\rightarrow$ $\mathbb{R}$) (h0 : $\forall$ t, f t = t
 - have eq1 : x^4 = 5 * x^2 - 6 := by
@@
 This shows the capability of choosing the correct step not only depends on the correctness of the thoughts but also at the confidence of the model at the time of generation.
-Example of mathd_algebra_313 where the intermediate step is incorrectly annotated.
+Example of Mathd Algebra 313 where the intermediate step is incorrectly annotated.
 The failure arises from the premature application of the simplification tactic field_simp immediately after the symbolic rewrite of $i$ as $v/z$, based on the hypothesis $v=i\cdot z$.
@@
 The appropriate use of field_simp should follow substitution of the concrete definitions for both $v$ and $z$, ensuring the resulting expression is a well-defined complex fraction amenable to symbolic manipulation and numeric normalization.
-Lean Proof for mathd_algebra_313
+Lean Proof for Mathd Algebra 313
 import Mathlib
@@
 The result highlights the non-trivial interaction between the encoded reasoning direction and the target proof environment, underscoring the potential of steering vectors as cross-system inductive biases.
-Problem Statement for Putnam_1988_B1
+Problem Statement for Putnam 1988 B1
 A composite number is a positive integer that can be expressed as a product (ab), where (a) and (b) are (not necessarily distinct) integers from the set ({2, 3, 4, dots}).
@@
 where (x, y, z) are positive integers.
-Putnam_1988_B1 proof in Coq
+Putnam 1988 B1 proof in Coq
 Require Import Arith Lia Ring.
@@
 exists x y, 1 < x /\ 1 < y /\ n = x * y.
-Lemma Putnam_1988_B1 n :
+Theorem Putnam_1988_B1 n :
 3 < n ->